# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v2

@@ B.9 Numerical Proof of Proposition 7
B.10 Formal Statement of Theorem 10 and Proof
B.10.1 Preliminaries of single-token prediction
@@
B.10.2 Formal statement of Theorem 10
@@
Recall that in DTSP task, we have $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)$, where $\psi(y),\psi(y,\omega),\textbf{r}_{\textup{sparse}}\in\mathbb{R}_{d}$, and $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, $k\ll d$. We further assume $B,L\in\mathbb{R}_{+}$, $\Theta_{B}:=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}},e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}$, $\|\psi(y)\|_{2}\leq L$, and $\psi(y,\omega)=\psi(\omega)+(\textbf{r}_{\textup{dense}}^{\top}\psi(y))e_{1}$.
@@
For DTSP task, we first collect a single-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{n}$, and ask human annotators to label these pairs. Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$, where $y_{w}^{(i)}=\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)}$ and $y_{l}^{(i)}=\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}$. And we further assume that the Gram matrix $\Sigma_{\mathcal{D}}:=\frac{1}{n}\sum_{i=1}^{n}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))^{\top}$ is non-singular, and $\text{tr}(\Sigma_{\mathcal{D}}^{-1})=\mathcal{O}(d)$.
@@
Under token-level linear parameterization and Assumptions 11 and 12, there exists an environment for DTSP task, s.t. by estimating from a preference dataset $\mathcal{D}$ with $n$ samples under $\theta_{1}=e_{1}$ constraint, the estimation error of the reward model $\hat{\theta}_{r}$ can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator:
@@
B.10.3 Proof of Theorem 13
Let $\pi_{{\mathsf{ref}}}(\cdot|y)$ be identical for all $y$, then we have
$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})=\textbf{r}_{\textup{dense}}^{\top}\psi(y)+C_{5}\;,$
for $\forall y$, where $C_{5}\in\mathbb{R}$ is an offset.
Recall that:
$(\theta_{r,0}^{\star})^{\top}\psi(y)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+C_{3}\;,$
$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(y,\omega)_{1})+C_{4}\;,$
@@ B.10.3 Proof of Theorem 13
$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)})-r_{\theta}(\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
@@
$\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\textbf{r}_{\textup{sparse}})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
@@
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\theta_{r,\textup{rel}\ell_{1},0}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}-\theta_{r,\textup{rel}\ell_{1},0})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
@@
$\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))=\log\sigma(\beta(\theta_{p,0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
then Lemma 1 implies that $\exists N=\mathcal{O}(d)$, for $\forall n\geq N$ and any estimator $\hat{\theta}_{p}$, we have
$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
@@
$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))-(\hat{\theta}_{p}+e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))\right]^{2}\;.$
@@
B.10.4 Proof of Lemma 4
@@
B.11 Omitted Calculations
Calculation of token-level structure of the optimal solution for DPO. As motivated by Rafailov et al. [2024], we show the token-level structure of the optimal solution for DPO as:
@@ B.11 Omitted Calculations
$\pi^{\star}(y_{t}|y_{0\ldots t-1})$ $=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})-q^{\star}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,$
[…]
To prove this, we define a $q^{\prime}$ function as:
For the initial token, by definition we have:
And then for a $y$ with $y_{N}$ as the terminal token, we have:
$=-\beta\log Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
thus
$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (10)
Then by definition:
we have:
and thus

### v3

@@ B.9 Numerical Proof of Proposition 7
B.10 Calculation of Token-level Structure of the Optimal Solution for DPO.
As motivated by Rafailov et al. [2024], we show the token-level structure of the optimal solution for DPO as:
$\pi^{\star}(y_{t}|y_{0\ldots t-1})$ $=\pi_{{\mathsf{ref}}}(y_{t}|y_{0\ldots t-1})\exp\left(\frac{q^{\star}(y_{t}|y_{0\ldots t-1})-q^{\star}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)\;,\;(t\neq 0)$
[…]
To prove this, we define a $q^{\prime}$ function as:
For the initial token, by definition we have:
And then for a $y$ with $y_{N}$ as the terminal token, we have:
$=-\beta\log Z+q^{\prime}(y_{N}|y_{0\ldots N-1})\;.$
thus
$q^{\prime}(y_{N}|y_{0\ldots N-1})=r^{\star}(y)\;.$ (11)
Then by definition:
we have:
and thus
$\sum_{s}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp\left(\frac{q^{\prime}(s|y_{0\ldots t-1})-q^{\prime}(y_{t-1}|y_{0\ldots t-2})}{\beta}\right)=1\;,$
which yields:
$q^{\prime}(y_{t-1}|y_{0\ldots t-2})=\beta\log\sum_{s\in\mathcal{V}}\pi_{{\mathsf{ref}}}(s|y_{0\ldots t-1})\exp(q^{\prime}(s|y_{0\ldots t-1})/\beta)\;.$ (13)
Combining Equations 10, 12, 11 and 13, we show that $q^{\star}$ exists and is equivalent to $q^{\prime}$.
Furthermore, if we assume $r^{\star}(y_{0\ldots t})$ to be decomposed as $\sum_{i=0}^{t}r^{\star}(y_{i}|y_{0\ldots i-1})$, we can define a soft value function $v^{\star}$:
[…]
Then samely, we have
$\pi^{\star}(y_{t}|y_{0,\ldots t-1})\propto\pi_{{\mathsf{ref}}}(y_{t}|y_{0,\ldots t-1})\exp\left(\frac{r^{\star}(y_{t}|y_{0\ldots t-1})+v^{\star}(y_{t}|y_{0\ldots t-1})}{\beta}\right)\;.$
Proof is the same as what we do on $q^{\star}$ so we omit it.
B.11 Formal Statement of Theorems 10 and 11 and Proofs
B.11.1 Preliminaries of single-token prediction
@@
B.11.2 Formal statement of Theorem 10
@@
Recall that in DTSP task, we have $r^{\star}(a,b)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\beta e_{1}^{\top}\psi(a,b)$, where $a,b\in\mathcal{V}$, $\psi(a),\psi(a,b),\textbf{r}_{\textup{sparse}}\in\mathbb{R}_{d}$, and $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, $k\ll d$. We further assume $B,L\in\mathbb{R}_{+}$, $\Theta_{B}:=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}},e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}$, $\|\psi(a)\|_{2}\leq L$, and $\psi(a,b)=\psi(b)+(\textbf{r}_{\textup{dense}}^{\top}\psi(a))e_{1}$.
@@
For DTSP task, we first collect a single-token dataset $\mathcal{D}^{\dagger}=\{a_{1}^{(i)},a_{2}^{(i)}\}_{i=1}^{n}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{a_{1}^{(i)}a_{1}^{(i)},a_{2}^{(i)}a_{2}^{(i)}\}_{i=1}^{n}$, and ask human annotators to label these pairs. Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$, where $y_{w}^{(i)}=a_{1}^{(i)}a_{1}^{(i)}$ and $y_{l}^{(i)}=a_{2}^{(i)}a_{2}^{(i)}$ w.p. $\sigma(r^{\star}(a_{1}^{(i)},a_{1}^{(i)})-r^{\star}(a_{2}^{(i)},a_{2}^{(i)}))$. And we further assume that the Gram matrix $\Sigma_{\mathcal{D}}:=\frac{1}{n}\sum_{i=1}^{n}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))^{\top}$ is non-singular, $\text{tr}(\Sigma_{\mathcal{D}}^{-1})=\mathcal{O}(d)$, and $n\geq c_{1}\text{tr}(\Sigma_{\mathcal{D}}^{-1})$, where $c_{1}$ is the constant in Lemma 1.
@@
Under token-level linear parameterization and Assumptions 12 and 13, there exists an environment for DTSP task, s.t. by estimating from a preference dataset $\mathcal{D}$ with $n$ samples under $\theta_{1}=e_{1}$ constraint, the estimation error of the reward model $\hat{\theta}_{r}$ can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator:
@@
B.11.3 Proof of Theorem 14
Let $\pi_{{\mathsf{ref}}}(\cdot|a)$ be identical for all $a$, then we have
$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|a)}{\mathbb{E}}\exp(\psi(a,b)_{1})=\textbf{r}_{\textup{dense}}^{\top}\psi(a)+C_{5}\;,$
for $\forall a\in\mathcal{V}$, where $C_{5}\in\mathbb{R}$ is a constant offset.
Note that for the second token, $\theta_{r}^{\star}$ and $\theta_{p}^{\star}$ share the same optimal solution:
$(\theta_{r,1}^{\star})^{\top}\psi(a,b)=e_{1}^{\top}\psi(a,b)+C_{1}\;,$
$(\theta_{p,1}^{\star})^{\top}\psi(a,b)=e_{1}^{\top}\psi(a,b)+C_{2}\;,$
where $C_{1},C_{2}\in\mathbb{R}$ are constant offsets when fixing $a$. And for the first token $a$, there is a distinction:
$(\theta_{r,0}^{\star})^{\top}\psi(a)$ $=\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+C_{3}\;,$
$(\theta_{p,0}^{\star})^{\top}\psi(a)$ $=\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|a)}{\mathbb{E}}\exp(r^{\star}(a,b)/\beta)+C_{4}$
$=\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\log\underset{w\sim\pi_{{\mathsf{ref}}}(\cdot|a)}{\mathbb{E}}\exp(\psi(a,b)_{1})+C_{4}$
$=\textbf{r}_{\textup{sparse}}^{\top}\psi(a)+\textbf{r}_{\textup{dense}}^{\top}\psi(a)+C_{4}+C_{5}\;,$
where $C_{3},C_{4}$ are constant offsets.
@@ B.11.3 Proof of Theorem 14
$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(a_{w}^{(i)}a_{w}^{(i)})-r_{\theta}(a_{l}^{(i)}a_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
$\implies\hat{\theta}_{r,\textup{rel}\ell_{1},0}$ $\in\underset{\theta_{0}+\tau_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\beta(\theta_{0}+\tau_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))+\gamma\|\theta_{0}+\tau_{1}-\tau_{1}\|_{1}\;,$
@@
$\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{r,\textup{rel}\ell_{1},0}-\textbf{r}_{\textup{sparse}})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
@@
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))-(\hat{\theta}_{r,\textup{rel}\ell_{1},0}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,\textup{rel}\ell_{1},0})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}$
@@
$\log\sigma(\hat{r}_{\theta_{p}}(a_{w}^{(i)}a_{w}^{(i)})-\hat{r}_{\theta_{p}}(a_{l}^{(i)}a_{l}^{(i)}))=\log\sigma(\beta(\theta_{p,0}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)})))\;,$
then Lemma 1 implies that for any estimator $\hat{\theta}_{p}$, we have
$\underset{e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\sup}\frac{1}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
@@
$\quad\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(\hat{r}_{\hat{\theta}_{p}}(y_{w}^{(i)})-\hat{r}_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\textbf{r}_{\textup{sparse}}+\textbf{r}_{\textup{dense}}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))-(\hat{\theta}_{p}+e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}$
$=\frac{\beta^{2}}{n}\sum_{i=1}^{n}\left[(\hat{\theta}_{p,0}+e_{1}-\textbf{r}_{\textup{sparse}}-\textbf{r}_{\textup{dense}}-e_{1})^{\top}(\psi(a_{w}^{(i)})-\psi(a_{l}^{(i)}))\right]^{2}\;.$
@@
$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(\hat{r}_{\hat{\theta}_{p}}(y_{w}^{(i)})-\hat{r}_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
B.11.4 Proof of Lemma 4
@@
B.11.5 Formal statement of Theorem 11 and proof
Lemma 6 (Lemma J.5 of Nika et al. [2024]).
If the features $\psi(a)$ are sampled from a 0-mean distribution and span $\mathbb{R}^{d}$, then $\log\sum_{a}\exp(\theta^{\top}\psi(a))$ is $\kappa$-strongly convex w.r.t. $\theta\in\Theta_{B}$, where $\kappa$ is an $\mathcal{O}(1)$ constant determined by $\beta,B,L$ and $|\mathcal{V}|$.
Theorem 15 (Formal sub-optimality separation theorem).
In the same setting as Theorem 14, there exists an environment for DTSP task, s.t. the sub-optimality of the RLHF policy model $\pi_{\textup{RLHF}}=\underset{\pi\in\Pi}{\operatorname{argmax}}\;V_{r_{\hat{\theta}_{r}}}^{\pi}$ can be reduced to $\mathcal{O}\left(\sqrt[4]{\frac{k\log d+k\log(1/\delta)}{n}}\cdot\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)$, i.e. w.p. $1-\delta$,
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{RLHF}}}$ $=\mathcal{O}\left(\sqrt[4]{\frac{k\log d+k\log(1/\delta)}{n}}\cdot\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)\;,$
while the sub-optimality of the DPO policy model $\pi_{\textup{DPO}}=\pi_{\hat{\theta}_{p}}$ is lower bounded:
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{DPO}}}$ $=\Omega\left(\frac{d}{n}\cdot\frac{1}{\|\Sigma_{\mathcal{D}}\|_{2}}\right)\;.$
Proof.
The proof follows the ideas of Theorem 3.2 of Zhu et al. [2023] and Theorem 4.2 of Nika et al. [2024], with appropriate adaptations to our setting.
@@ B.11.5 Formal statement of Theorem 11 and proof
$V_{r^{\star}}^{\pi^{\star}}-V_{r^{\star}}^{\pi_{\textup{RLHF}}}$ $\leq\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},b_{1}\sim\pi^{\star}(\cdot|a_{1}),\\ a_{2}\sim\pi_{\textup{RLHF}},b_{2}\sim\pi_{\textup{RLHF}}(\cdot|a_{2})\end{subarray}}{\mathbb{E}}\left[\left(r^{\star}(a_{1},b_{1})-r^{\star}(a_{2},b_{2})\right)-\left(\beta\log\frac{\pi_{\text{RLHF}}(a_{1},b_{1})}{\pi_{{\mathsf{ref}}}(a_{1},b_{1})}-\beta\log\frac{\pi_{\text{RLHF}}(a_{2},b_{2})}{\pi_{{\mathsf{ref}}}(a_{2},b_{2})}\right)\right]$
$=\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},b_{1}\sim\pi^{\star}(\cdot|a_{1}),\\ a_{2}\sim\pi_{\textup{RLHF}},b_{2}\sim\pi_{\textup{RLHF}}(\cdot|a_{2})\end{subarray}}{\mathbb{E}}\left[\left(r^{\star}(a_{1},b_{1})-r^{\star}(a_{2},b_{2})\right)-\left(r_{\hat{\theta}_{r}}(a_{1},b_{1})-r_{\hat{\theta}_{r}}(a_{2},b_{2})\right)\right]$
$=\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},\\ a_{2}\sim\pi_{\textup{RLHF}}\end{subarray}}{\mathbb{E}}\left[\beta(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0})^{\top}(\psi(a_{1})-\psi(a_{2}))\right]$
$=\beta(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0})^{\top}\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},\\ a_{2}\sim\pi_{\textup{RLHF}}\end{subarray}}{\mathbb{E}}(\psi(a_{1})-\psi(a_{2}))$
$\leq\beta\|\Sigma_{\mathcal{D}}^{1/2}(\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0})\|_{2}\|\Sigma_{\mathcal{D}}^{-1/2}\underset{\begin{subarray}{c}a_{1}\sim\pi^{\star},\\ a_{2}\sim\pi_{\textup{RLHF}}\end{subarray}}{\mathbb{E}}(\psi(a_{1})-\psi(a_{2}))\|_{2}$
$=\beta\|\textbf{r}_{\textup{sparse}}-\hat{\theta}_{r,0}\|_{\Sigma_{\mathcal{D}}}\cdot\mathcal{O}\left(\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)$
$=\mathcal{O}\left(\sqrt[4]{\frac{k\log d+k\log(1/\delta)}{n}}\cdot\left\|\Sigma_{\mathcal{D}}^{-1/2}\right\|_{2}\right)\;.$
The first inequality comes from performance difference lemma (see Section B.12); the second equality comes from the observation that all $r_{\theta_{r}}$ with $\theta_{r,1}=e_{1}$ can be fitted by the log-linear policy model; the third and fourth equalities come from simple calculations under our setting; the fifth inequality comes from Cauchy-Schwarz inequality; the sixth equality comes from the fact that $\psi(a)$ is bounded; and the last equation comes from Theorem 14.
Since the optimal policy satisfies $\pi^{\star}(a,b)=\pi_{{\mathsf{ref}}}(a,b)\exp(r(a,b)/\beta)/Z$, we have: